# Intermediate mass-ratio inspirals in a dense dark-matter environment: Effects of the initial dark-matter distribution

### v1

@@ B.2 Nonzero angular-momentum cutoff
$I_{\mathcal{E}}(\Delta\mathcal{E})=\begin{cases}4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{1},m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})<\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}<\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2}),\ \mbox{ and }\\ &\phantom{for}\ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+2\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{1},m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+2\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ \sqrt{2[\mathcal{E}_{\mathrm{max}}(r_{2})-\mathcal{E}]}\,(\alpha_{2}-\alpha_{1})&\mbox{for}\ \ r_{2}=r_{\mathcal{J}},\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{2},m\right)&\mbox{for}\ \ r_{\mathcal{J}}(1-8\sqrt{q}/3)\leq r_{2}<r_{\mathcal{J}},\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(8q/3)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}}\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 0&\mbox{otherwise}.\end{cases}$ (77)
@@
$I_{\mathcal{E}}(\Delta\mathcal{E})=\begin{cases}4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{1},m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})<\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}<\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{1},m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ \sqrt{2[\mathcal{E}_{\mathrm{max}}(r_{2})-\mathcal{E}]}\,(\alpha_{2}-\alpha_{1})&\mbox{for}\ \ r_{2}=r_{\mathcal{J}},\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}(r_{2}),\\ 4\sqrt{\dfrac{|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{2},m\right)-E\left(\phi_{1},m\right)]&\mbox{for}\ \ r_{\mathcal{J}}(1-4\sqrt{q})\leq r_{2}<r_{\mathcal{J}},\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 0&\mbox{otherwise}.\end{cases}$ (78)

### v2

@@ B.2 Nonzero angular-momentum cutoff
$I_{\mathcal{E}}(\Delta\mathcal{E})=\begin{cases}4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{1},m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})<\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+2q),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}<\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2}),\ \mbox{ and }\\ &\phantom{for}\ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+2\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{1},m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+2\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(2q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ \sqrt{2[\mathcal{E}_{\mathrm{max}}(r_{2})-\mathcal{E}]}\,(\alpha_{2}-\alpha_{1})&\mbox{for}\ \ r_{2}=r_{\mathcal{J}},\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{2},m\right)&\mbox{for}\ \ r_{\mathcal{J}}(1-8\sqrt{q}/3)\leq r_{2}<r_{\mathcal{J}},\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(8q/3)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}}\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 0&\mbox{otherwise}.\end{cases}$ (77)
@@
$I_{\mathcal{E}}(\Delta\mathcal{E})=\begin{cases}4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}E\left(\phi_{1},m\right)&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})<\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{2}\geq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\mathrm{min}(\Delta\mathcal{E}_{\mathrm{up}},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)}),\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}<\mathcal{E}_{\mathrm{cut}}^{(+b_{\star})}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{1},m\right)-E(\phi_{2},m)]&\mbox{for}\ \ r_{\mathcal{J}}<r_{2}\leq r_{\mathcal{J}}(1+4\sqrt{q}),\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ \sqrt{2[\mathcal{E}_{\mathrm{max}}(r_{2})-\mathcal{E}]}\,(\alpha_{2}-\alpha_{1})&\mbox{for}\ \ r_{2}=r_{\mathcal{J}},\ \ -v_{2}^{2}\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{cut}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathcal{E}_{\mathrm{max}}(r_{2}),\\ 4\sqrt{\dfrac{b_{\star}|\mathcal{E}^{\prime}_{\mathrm{max}}(r_{2})|}{m}}[E\left(\phi_{2},m\right)-E\left(\phi_{1},m\right)]&\mbox{for}\ \ r_{\mathcal{J}}(1-4\sqrt{q})\leq r_{2}<r_{\mathcal{J}},\ \ \mathrm{max}(-v_{2}^{2},\Delta\mathcal{E}_{\mathcal{J}}^{(4q)})\leq\Delta\mathcal{E}\leq\Delta\mathcal{E}_{\mathrm{up}},\\ &\phantom{for}\mathcal{E}_{\mathrm{max}}^{(-b_{\star})}(r_{2})\leq\mathcal{E}\leq\mathrm{min}[\mathcal{E}_{\mathrm{max}}^{(+b_{\star})}(r_{2}),\mathcal{E}_{\mathrm{max}}(r_{\mathcal{J}})],\\ 0&\mbox{otherwise}.\end{cases}$ (78)